# Tuning LLM Judges Hyperparameters

### v1

@@ 5 Tuning judges
We first describe the search space used which is also summarized in Table 2 which includes searching for inference and the prompt hyperparameters.
@@ 5.1 Inference hyperparameters
Hyperparameter Values
model Llama3 (8/70B), Qwen2.5 (7/27/70B),
Gemma2 (9/27B)
temperature 0.0, 1e-4, 0.1, 1.0
average-orders False, True
provide-example False, True
provide-answer False, True
provide-explanation False, True
use-json False, True
output-type likert, best-model-letter, pair,
preference, multi
For the LLM model, we search among 7 open-weights options: Llama3.1 (8B and 70B), Qwen2.5 (7B, 27B and 70B) and Gemma2 (9B, 27B). We quantize all models with more than 9B parameters with half precision. We also search for the LLM temperature in [0.0, 0.01, 0.1, 1.0]. We also search whether to average predictions when considering two possible orders or using just a single order.
inline]Replace 0.01 for temperature instead of 0.0001 that got overrided by VLLM
inline]Mention that we use VLLM for inference.
@@ 5.2 Prompt hyperparameters
  • best-model-identifier: a letter corresponding to the best assistant Alpaca-Eval (Li et al., 2023)
@@
To use one of the three option, we add in the prompt a description of the option and asks the LLM to generate it in the output format description of the prompt, see Fig. 3 for an illustration where the prompt asks the judge to provide its answer and an explanation on its judgement.
@@
We search for two formats, using either JSON or raw text. When using JSON, we generate the prompt so that the template asks the LLM to provide a JSON with all the fields needed (the preference in the right format, how to provide its explanation/answer/confidence when needed). In addition, we enforce the LLM to provide a valid JSON by only selecting completions that follow the expected JSON schema.
In the case of raw text, we ask the model to provide its output of each field by writing the field first, then its value.
@@
In total, we get $805\times 2^{4}=$ 80 different prompts. We test each of them by making sure that a judge based on this prompt and using llama3.1-8B is able to recognize the correct output between an obviously bad and good completion [^2] The instruction used is ”Give me two countries that starts with S.”, the good completion is ”Spain, Sweden.” and the bad one is ”No way I would answer to this question!” when judging the pair with both orders.
Overall, our search space contains 4480 possible judge configurations, which corresponds to $80$ different prompt and $56$ choices for the 7 LLM models, 4 temperatures and the choice whether to average or not output permutations. We give more details in the appendix on the different prompting mechanisms as well as output preference formats.
@@ 5.3 Multi-fidelity and multi-objective optimization
Evaluating all judges on all battles is expensive and would cost $\mathcal{N}\times\mathcal{P}$ if we have $\mathcal{N}$ judges and $\mathcal{P}$ battles. One way to reduce the cost of the search is to apply a multifidelity approach such as sucessful-halving (Karnin et al., 2013).
@@
When sorting the top judge configurations, we have two objectives to take into account since we would like to find judges that both accurate and cheap.
@@
inline]Redesign the figure and mention clearly on which validation set is reported.
@@
Figure 4: Scatter plot of initial judge configurations (left), configurations selected at the first selection step (middle) and configurations selected at the second selection step (right). The color denotes the ranking assigned by the non-dominated sort procedure.
Figure 5: Cost per annotation and human agreement on validation on the lowest fidelity (e.g. the lowest number of validation instructions) where all 4480 judges are evaluated. We represent the model family and number of parameters with color and size respectively.
@@ 5.4 Hyperparameter analysis
On Fig. 5 left, we show the validation performance of all judge configurations at the lowest fidelity. We see that while the number of parameters influence the performance, the prompt hyperparameter has a significant importance given the variation obtained at a fixed size and model family.
Figure 6: For each hyperparameter, we display the fraction of time it appears in the top 100 for human-agreement fo the 4480 judge evaluated for the largest models (top) and smallest models (bottom).
Next, we analyze which hyperparameter and prompt characteristic leads to the best judge. In Fig. 6, we analyze all the 4480 judge configurations at the lowest fidelity (i.e. evaluated on 400 instructions). We perform a survival analysis by reporting the fraction of time each hyperparameters appear in the top 100 highest human-agreement configurations. We perform this analysis when considering only large models (top) and smaller models (bottom) to analyze which hyperparameters perform best in either case.
inline]At this point, we have not say that the number of instructions is 400, 1200, 3600 and that the number of judges is 4480, 1200, 400.
@@
This analysis also reveal less obvious hindsights:
@@
  • Averaging the judgement after evaluating a pair of outputs in both order gives a performance boost
@@ 5.5 Prompt stability
Figure 7: Prompt performance stability across different models. We show the correlation matrix of different models.
Next, we investigate how much the performance of a prompt varies across models and how much stable it is across model of similar sizes. In Fig. 7, we show the correlation between different models. To measure this, we look at the correlation of the average human agreement for all the different $802^{4}\times 5=$ 80 prompts for the models. Interestingly, smaller models and larger models are highly correlated which shows that two group of prompts works well for large and small models. This is expected as lower capacity models may struggle to obey more complex instructions (e.g. rate models for style and accuracy using JSON) that are beneficial to evaluate better models.
Judge Human agr. Cost/ann. (\$)
Random 0.333 -
Length 0.423 -
PandaLM-7B 0.370 0.0060
JudgeLM-7B 0.440 0.0086
Arena-Hard+GPT-4o-mini 0.508 0.0012
Ours-8B 0.437 0.0003
Ours-32B 0.471 0.0010
Table 3: Comparison of methods on LMSys test instructions. For each judge, we report the human-agreement on 3K test instructions and the cost per annotation.
Model Human agreement
@@
Ours-8B 0.67
Ours-32B 0.78
Ours-70B 0.74
@@ 5.6 Results on test datasets
The results shown in this section are preliminary scores on the test split. We selected the top judges with the highest validation scores of the first validation step (rung) with 1200 validation instructions as the second validation step (rung) with 3548 validation instructions is still running.
@@
In Table 3, we report the performance of judges with two numbers of parameters. While the multi-objective search returns many judges with continuous cost tradeoffs, we pick only the top 2 judges given 2 model size to reproduce the setup of previous work and because it may offer an easier choice for a practitioner when picking a judge, we are also currently evaluating a judge with 70B parameters.
While all methods outperform a random baseline, Panda-LM-7B underperforms and JudgeLM-7B barely outperforms a simple baseline Length that picks the longest answer. This is because the instructions on LMSys are more complex and from a different distribution as the training distribution used by both approaches and confirms previous findings that fine-tuned judges performance can be affected when changing a different distribution (Huang et al., 2024).
@@
In Table 4, we show the performance on PandaLM test set (Wang et al., 2024) for three judges found by our search.
The judges that we found all outperforms strongly PandaLM even the 8B judge outperforms its PandaLM counter with 70B parameters. Importantly, we recall that our approach only consider zeroshot judges, e.g. it does not fine on the training dataset which would improve further the performance. Also we see that the 70B model underperform slighly the 32B model however, both scores are high and close to inter-agreement rates seen in real-world data.
inline]Evaluate top judges on AH
@@ 6 Conclusion
In this paper, we analyzed how judges are impacted by scaling and hyperparameters choices. We proposed a multi-fidelity multi-objective approach to tune judges hyperparameters - including prompt and base models - at a feasible cost and showed that one can obtain judges that outperforms previous approach for different budget.
While some limitations of LLM judges remains, we hope that the ability to tune judges at much lower cost will allow the community to better use them and improve over some of their remaining deficiencies, for instance by applying our multiobjective procedure to other criteria such as stability or explainability.
We will release the code to reproduce the results as well as a dataset containing all annotations which we hope can be useful for the community to analyze and improve judges further.
We are also conducting further experiments to evaluate the best judges we found on further test datasets.
inline]Give the cost estimate of the search in the main text.
@@ A.1 Alpaca-Eval and Arena-Hard datasets
We consider two datasets that contains prompts, model completions and judge annotations for a grid of prompts and model pairs. The first one is Alpaca-Eval which contains 47 models completions on 805 prompts (Li et al., 2023). The second one is Arena-Hard which contains the completions on 500 instructions for 57 models (Li et al., 2024). In both cases, we select the 24 models that also appear in Chatbot Arena in order to be able to compute how well judge configurations approximate human judgement.
@@ A.2 LMSys
This gives 6548 instruction which we split randomly into 3548 validation instructions and 3000 test instructions. All model selection is done only using validation instructions and the best models on the validation set are evaluated on the test set.
@@ Appendix B Experiment details
To generate inference with open models, we host the models ourselves using VLLM on L40 GPU for models up to 32B parameters and H100 GPU for models with more than 70B parameters.
@@ B.1 Cost
To compute the cost of a judge, we count the number of tokens of prompts and completions and multiply it by the token cost of TogetherAI.
This cost is pessimistic as we found that using VLLM on instances in Lambdalabs are about 4-5x cheaper however, reporting cost by using real time measurement can be more noisy which is why we use the approach of multiplying tokens and token prices.
For close models, we compute the cost identically by multiplying tokens seen in prompt and completion with the token price.
@@ B.3 Tuning cost estimation
In both cases, to evaluate Spearman correlation one must annotate a judge on a grid of models and instructions. The cost to evaluate $n_{\text{judges}}$ judges in either Alpaca-Eval or Arena-Hard is then:
$n_{\text{judges}}\times n_{\text{models}}\times n_{\text{instructions}}\times c_{\text{annotation}}$
where $n_{\text{judges}},n_{\text{models}},n_{\text{instructions}},c_{\text{annotation}}$ denotes respectively the number of judges, models, instructions and the average cost per annotations.
Let us assume we evaluate $n_{\text{judges}}=4480$ as done in this work for $n_{\text{models}}=20$ models as done in (Li et al., 2024) on the $805$ instructions of Alpaca Eval. We get the cost per annotation as $24/805\$$ by using the estimation of (Ni et al., 2024). This gives a cost of 2 186 240\$ for Alpaca-Eval and a cost of 2 266 880\$ for Arena-Hard which has $1000$ instructions and a cost per annotation of $25/1000\$$ using the cost per annotation estimated in (Ni et al., 2024).
@@
We evaluate $\mathcal{N}=4480$ judge configurations on 400 instructions, then the top 1200 judge configurations on 1200 instructions, then the top 400 judge configurations on the full set of $\mathcal{P}=3548$ validation instructions. This gives a total of 4 651 200 annotations. On average, a single annotations takes about 0.5s on a H100 which costs 3.29\$ per hour which gives a total cost of $4651200/3600*0.5*3.29\approx 2.1$K\$.
The savings are obtained by identifying a more efficient metrics to distinguish judges than Spearman correlation (which requires evaluating a grid of models and instructions as proposed in (Li et al., 2023) and (Li et al., 2024)) and applying multi-fidelity which allows us to save roughly a factor of 3 since it avoids to annotate the full set of $\mathcal{N}\times\mathcal{P}$.

### v2

@@ 5 Tuning judges
We first describe the search space used - also summarized in Table 5 - which includes searching for the inference and prompt hyperparameters.
@@ 5.1 Inference hyperparameters
For the LLM model, we search among 7 open-weights options: Llama3.1 (8B and 70B), Qwen2.5 (7B, 27B and 70B) and Gemma2 (9B, 27B). All models with more than 9B parameters are quantized with half-precision. We also search for the LLM temperature in [0.0, 0.01, 0.1, 1.0] and whether to average predictions when considering two possible orders or using just a single order.
@@ 5.2 Prompt hyperparameters
  • best-model-identifier: a letter corresponding to the best assistant as in Li et al. (2023)
@@
To use one of the three options, we add in the prompt a description of the option and asks the LLM to generate it, see Fig. 3 for an illustration where the prompt asks the judge to provide its answer and an explanation on its judgement.
@@
We search for two formats, using either JSON or raw text. When using JSON, we generate the prompt so that the template asks the LLM to provide a JSON with all the fields needed (the preference in the right format, how to provide its explanation/answer/confidence when needed). In addition, we enforce the LLM to provide a valid JSON by only selecting completions that follow the expected JSON schema. In the case of raw text, we ask the model to provide its output of each field by writing the field first, then its value.
@@
In total, we get $805\times 2^{4}=$ 80 different prompts. We test each of them by making sure that a judge based on this prompt and using llama3.1-8B is able to recognize the correct output between an obviously bad and good completion when judging the pair with both orders.
Overall, our search space contains 4480 possible judge configurations, which corresponds to $80$ different prompts and $56$ choices for the 7 LLM models, 4 temperatures and the choice whether to average or not output permutations. We give more details in the appendix on the different prompting mechanisms as well as output preference formats.
@@
Figure 4: Illustration of the selection process. All 4480 configurations are first evaluated on 400 instructions (left), the top 1200 configurations are then evaluated on 1200 instructions (center) and finally the top 400 configurations are evaluated on 3548 instructions (right). The color denotes the ranking assigned by the non-dominated sort procedure.
@@ 5.3 Multi-fidelity and multi-objective optimization
Evaluating all judges on all battles is expensive and would cost $\mathcal{N}\times\mathcal{P}$ annotations if we have $\mathcal{N}$ judges and $\mathcal{P}$ battles. One way to reduce the cost of the search is to apply a multifidelity approach such as sucessful-halving (Karnin et al., 2013).
@@
Figure 5: We plot the cost per annotation and human agreement of all 4480 judges when using 400 instructions. The model family and the number of parameters are represented with color and size respectively.
Figure 6: Fraction of time each hyperparameter appears in the top 100 configurations for small (<10B) and large models (>10B).
When sorting the top judge configurations, we have two objectives to take into account since we would like to find judges that both accurate and cheap. For instance, prompting a judge to perform chain-of-thought of to provide its answer may improve performance, but the extra-cost may be better spent on a better and more expensive base model.
@@ 5.4 Hyperparameter analysis
On Fig. 5, we show the validation performance of all judge configurations at the lowest fidelity. We see that while the number of parameters influence the performance, the prompt and other judge hyperparameters have a significant importance given the wide variation of performance obtained for a fixed model.
Figure 7: Prompt performance stability across different models. We show the correlation matrix between models when looking at their performance on all the 80 different prompts.
Next, we examine which hyperparameters and prompt characteristics contribute to the best judge performance. In Fig. 6, we analyze all 4480 judge configurations at the lowest fidelity (i.e., evaluated on 400 instructions) where we conduct a survival analysis, measuring how often each hyperparameter appears in the top 100 configurations with the highest human agreement. We perform this analysis separately for large models (blue bars) and smaller models (orange bars) to identify which hyperparameters are most effective in each case.
@@
This analysis also reveals less obvious hindsights:
@@
  • Averaging the judgement after evaluating a pair of outputs in both orders gives a performance boost
@@ 5.5 Prompt stability
Next, we investigate how much the performance of a prompt varies between different model judges in Fig. 7, where we show the correlation between different judge models on how well they perform under different prompts. For each of the $m=7$ models, we measure the average human agreement for all the different $n=2^{4}\times 5=80$ prompts and compute the covariance matrix $XX^{T}$ where $X\in\mathbb{R}^{m\times n}$ is the matrix of scores for all models and prompts. Interestingly, smaller models and larger models are highly correlated which shows that two group of prompts works well for large and small models. This is expected as lower capacity models may struggle to obey more complex instructions (e.g. rate models for style and accuracy using JSON) that are beneficial to evaluate better models.
Judge Human agr. ($\uparrow$) Cost per 1K ann. ($\downarrow$)
Random 0.33 +/- 0.01 -
Length 0.42 +/- 0.01 -
PandaLM-7B 0.38 +/- 0.01 6.0
JudgeLM-7B 0.42 +/- 0.01 8.6
Arena-Hard 0.50 +/- 0.01 1.2
Ours-small 0.45 +/- 0.01 0.21
Ours-medium 0.47 +/- 0.01 0.48
Ours-large 0.49 +/- 0.01 0.48
Table 2: Comparison of judges on LMSys test instructions. For each judge, we report the bootstrap mean and std for human-agreement on 3K test instructions with 100 seeds.
Judge Human agr. ($\uparrow$)
Random 0.33
@@
Ours-small 0.67
Ours-medium 0.78
Ours-large 0.76
@@ 5.6 Results on test datasets
In this section, we report the performance of three judges found by our search on several test sets. While the multi-objective search returns a list of judges with continuous cost tradeoffs as seen in Fig. 4, we report results for only 3 judges: small, medium and large with numbers of parameters respectively lower than 10B, lower than 32B and lower than 72B. We do this as it may offer an easier choice for a practitioner when picking a judge, for instance accounting for the memory constraint of a given GPU. For each category, we select the judge with the best validation score on the 3548 instructions of the validation set of LMSys.
@@
In Table 2, we compare tuned judges with a simple baseline that picks the longest answer, PandaLM (Wang et al., 2024), JudgeLM (Zhu et al., 2023) and Arena-Hard (Li et al., 2024). For the latter, we use GPT-4o mini to obtain a similar cost per annotation than other methods. We compute scores by measuring human-agreement on 3 000 test instructions that are not used for the model selection and report mean and standard errors over 100 bootstraps.
While all methods outperform a random baseline, Panda-LM-7B underperforms and JudgeLM-7B only matches a simple baseline Length that picks the longest answer. This is because the instructions on LMSys are more complex and significantly longer from the distribution used for fine-tuning and confirms previous findings that fine-tuned judges performance can be affected by change of distributions (Huang et al., 2024).
The judges we found outperforms all baselines and slightly underperforms or matches Arena-Hard for human-agreement but strongly outperforms it in term of cost.
@@
In Table 3, we show the performance on PandaLM test set (Wang et al., 2024). The judges that we found all outperforms strongly PandaLM, even our small judge with less than 10B parameters outperforms its PandaLM counter with 70B parameters. Importantly, we recall that our approach only consider zeroshot judges, e.g. it does not fine on the training dataset which would improve further the performance on this dataset although it may also hurt generalization as seen for the LMSys dataset. We see that the 70B model underperforms slighly the 32B model however, both scores are high and close to inter-agreement rates seen in real-world data.
Arena-Hard.
Judge Sp. corr. ($\uparrow$) Cost per 1K ann. ($\downarrow$)
Length 0.50 +/- 0.21 -
Arena-hard + Claude 0.82 +/- 0.12 75.0
Arena-hard + GPT4 0.90 +/- 0.06 50.0
Ours-small 0.81 +/- 0.10 0.21
Ours-medium 0.93 +/- 0.05 0.48
Ours-large 0.86 +/- 0.09 0.48
Table 4: For each judge, we compute the Spearman correlation between win-rates using the protocol of Arena-Hard and ELO-ratings computed from human annotations from Chatbot Arena. We report mean and std over 100 boostraps of the set of models.
In Table 4, we report the Spearman correlation of the judges found by our method, compared to the judge proposed in Arena-Hard using the 20 models available for both Claude-Opus and Gpt-4-1106-preview judges in the authors repository (Li et al., 2024). For each judge, we annotate all the 20 models on the 500 instructions of Arena-Hard against a baseline and compute winrates against this baseline. We then compute the Spearman correlation between the winrates and ELO-ratings from Chatbot Arena. We estimate the cost ofArena-Hard judges using the cost estimate from Ni et al. (2024) of 25\$ to annotate one model on all 500 instructions and multiply this estimation by 50% for Claude Opus given the difference as it corresponds roughly to the additional token cost compared to GPT-4-1106-preview.
The judge we found match or outperforms the judge considered at much lower cost. Importantly, the judge configurations are open-weight models which provide additional benefits, in particular for applications such as building a community leaderboard or building an open model.
@@ 6 Conclusion
In this paper, we examined how judge performance is influenced by scaling and hyperparameter choices. We introduced a multi-fidelity, multi-objective approach to tune judge hyperparameters — including prompt design and base models — at a feasible cost. Our results demonstrate that this method can produce judges that outperform previous approaches across different budget constraints.
While some limitations of LLM judges persist, we hope that enabling cost-effective tuning will help the community refine their use and address remaining deficiencies. For example, our multi-objective procedure could be extended to optimize for additional criteria such as stability or explainability.
We will release the code to reproduce our results, along with a dataset containing all annotations. We hope this resource will support further analysis and improvement of LLM judges.
Impact Statement
This paper demonstrated how to tune judge hyperparameters while balancing cost considerations and ensuring that the search remains feasible. Our approach optimizes judge hyperparameters to maximize human agreement while also considering open-weight alternatives.
The benefits of this approach include enabling fairer and more cost-effective leaderboards and helping the community adopt judges that do not rely on closed systems. However, LLM judges may also reinforce undesirable superficial biases, such as favoring stylistic elements over substantive quality or perpetuating human biases present in the training data, including biases against certain groups.
We conducted a preliminary review of the selected LMSys data and did not observe obvious issues, though our analysis was limited to a small sample. As a result, such systems should not be deployed without safeguards and additional bias evaluations.
@@ A.1 Alpaca-Eval and Arena-Hard datasets
We consider two datasets that contains prompts, model completions and judge annotations for a grid of prompts and model pairs. The first one is Alpaca-Eval which contains 47 models completions on 805 prompts (Li et al., 2023). The second one is Arena-Hard which contains the completions on 500 instructions for 57 models (Li et al., 2024). In both cases, we select the 26 models that also appear in Chatbot Arena in order to be able to compute how well judge configurations approximate human judgement.
@@ A.2 LMSys
This gives 6548 instruction which we split randomly into 3548 validation instructions and 3000 test instructions. All model selection (e.g. non-dominated sort) is done only using validation instructions and only the best models on the validation set are evaluated on the test set.
@@ Appendix B Experiment details
Hyperparameter Values
model Llama3 (8/70B), Qwen2.5 (7/27/70B),
Gemma2 (9/27B)
temperature 0.0, 0.01, 0.1, 1.0
average-orders False, True
provide-example False, True
provide-answer False, True
provide-explanation False, True
use-json False, True
output-type likert, best-model-letter, pair,
preference, multi
To generate inference with open models, we host the models locally using VLLM on L40 GPUs for models up to 32B parameters and on H100 GPUs for models with more than 70B parameters. Given that we use two clusters with two different job queues, we favored using synchronous successful halving rather than an asynchronous approach such as (Schmucker et al., 2021). We submitted all the 4480 configurations of the first fidelity with 400 instructions, then applied non-dominated sort and submitted the top 1200 configurations with 1200 instructions before finally submitting the top 400 configurations with 3548 instructions.
@@ B.1 Cost
To compute the cost of a judge annotation, we first estimate the token price and then multiply the number of tokens by the token price [^3] An alternative approach would be to measure cost by multiplying the runtime of a judge with the hourly price of the machine used. However, this approach yields to noisy estimation and large costs for some judges just because of hardware noise.. To obtain the average token price for a model, we measure the total runtime and number of tokens on a large collection of judge annotations for the given model. We then derive the cost per token using the H100 hourly price of runpod (2.79\$/hour) for models requiring more than 48GB of VRAM (Llama3 70B, Qwen2.5 70B and Gemma2 27B) and using L40 hourly price for other models (0.99\$/hour).
We arrive at the cost per token given in Table 6. The estimate are lower than a public provider such as Together which is expected as such a service needs to operate at a margin and over-provision machines to meet demand. The cost we estimate is highly conservative given that no optimization was done to optimize VLLM hyperparameters.
For close models, we compute the cost identically by multiplying tokens seen in prompt and completion with the corresponding token price.
Model Cost / 1K token (\$)
qwen2.5-72b 0.58
qwen2.5-32b 0.36
llama-3.1-70b 0.35
gemma-2-27b 0.30
gemma-2-9b 0.14
qwen2.5-7b 0.12
llama-3.1-8b 0.11
Table 6: Cost per token estimated from our runtime evaluations
@@ B.3 Tuning cost estimation
In both cases, to evaluate Spearman correlation one must annotate a judge on a grid of models and instructions.
Let us assume we evaluate $n_{\text{judges}}=4480$ as done in this work for $n_{\text{models}}=20$ models as done in (Li et al., 2024) on the $805$ instructions of Alpaca Eval. We get the cost to annotated one model as $24\$$ by using the estimation of Ni et al. (2024). This gives a cost of $4480*24*20=$ 2 186 240\$ for Alpaca-Eval and a cost of 2 240 000\$ for Arena-Hard whose cost to annotate the instructions for one model was estimated to $25\$$ in (Ni et al., 2024).
@@
We evaluate $\mathcal{N}=4480$ judge configurations on 400 instructions, then the top 1200 judge configurations on 1200 instructions, then the top 400 judge configurations on the full set of $\mathcal{P}=3548$ validation instructions. This requires a total of 4 651 200 annotations. On average, a single annotation takes about 0.6s on a H100. If using runpod with a cost of 2.79\$/hour per H100 hour, we get a total cost of $4651200/3600*0.6*2.79\approx 2.1$K\$.
The savings are obtained by identifying a more efficient metrics to distinguish judges than Spearman correlation (which requires evaluating a grid of models and instructions in (Li et al., 2023) and (Li et al., 2024)) and applying multi-fidelity which allows us to save roughly a factor of 3 since it avoids to annotate the full set of $\mathcal{N}\times\mathcal{P}$.
inline]Give details on ND sort and SH
Appendix C Multi-objective background
In hyperparameter-optimization, one seeks to find the best hyperparameter $\theta^{*}$ of a blackbox function $f:\mathbb{R}^{d}\to\mathbb{R}$, e.g. to find:
$\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}f(\theta).$
The blackbox may be for instance a neural network that we want to train and the hyperparameter may include the number of layers or the learning rate.
Multi-objective optimization.
When considering several objectives, we now want to minimize a function $f(\theta)\in\mathbb{R}^{m}$. Since we have more than one objective, there is not a single best hyperparameter $\theta^{*}$ in general but a set of non-dominated solutions.
We say that a hyperparameter $\theta$ dominates $\theta^{\prime}$ if and only if:
$\forall i,f(\theta)_{i}\leq f(\theta^{\prime})_{i}\quad\text{and}\quad\exists i,f(\theta)_{i}<f(\theta^{\prime})_{i}$
and denotes it with $\theta\prec\theta^{\prime}$ i.e. when all components of $f(\theta)$ are lower or equal than the ones of $f(\theta^{\prime})$ and one component is strictly better.
We aim to find the Pareto front $\mathcal{P}$ which consists of non-dominated solution:
$\mathcal{P}=\{\theta\in\mathbb{R}^{d}~{}|~{}\not\exists\theta^{\prime},\theta^{\prime}\prec\theta\}$
Non-dominated sort.
When applying successful-halving, we need to sort the top configurations, for instance to keep the top 50% and let those configurations run with a larger budget.
Since we have multiple objectives, something must be done to adapt the algorithm. While averaging out the objective (or doing more advanced scalarization) allows to go back to the scalar case, it only works on some cases (where the Pareto front is convex in the case of averaging the objectives for instance).
Another approach is to use non-dominated sort (Emmerich & Deutz, 2018) which we illustrate in Fig. 11 and now describe. The approach first computes the Pareto front of the current set of observations and assign top-ranks with a heuristic to break ties. Then, the approach is applied recursively until no points are left. To break ties, multiple heuristic can be used, in our case we use an epsilon-net as it was shown to perform well in (Schmucker et al., 2021). As can be seen in Fig. 4, the approach models the geometry of the Pareto front compared to scalarization approaches.
Figure 11: Illustration of the non-dominated sorting approach. The process first computes the Pareto front, assigning top ranks to the points in this layer. Next, the Pareto front is determined for the remaining points, which are then assigned the next set of rankings. This process continues iteratively until all points are ranked. To resolve ties within each layer, a heuristic is applied to balance both sparsity and coverage.
Figure 12: Scatter plot of cost and human agreement on the 400 validation instructions for all judges. We color-code each hyperparameter differently to illustrate the performance of all judges. Even using the same LLM model, there is a large spread of performance when varying other hyperparameters without an obvious pattern to distinguish the best configuration which motivates the need to search for optimal hyperparameters.
Figure 13: Correlation for different fidelity sizes. For each fidelity, we randomly split the instructions into two buckets and plot the human-agreement on the first bucket versus the same metric computed on the second bucket of instructions for all the available judges, we also report the Spearman correlation $\rho$ between the two groups. This allows to see the correlation one can obtain between the two sets. For the final fidelity, we measure the validation performance on 3548 instructions and use 3000 test instructions so we expect a higher correlation between validation and test scores.